OpenStudy (anonymous):

The function H described by H(x) = 2.75x + 71.48 can be used to predict the height, in centimeters, of a woman whose humerus (the bone from the elbow to the shoulder) is x cm long. Predict the height of a woman whose humerus is 35 cm long. The predicted height of a woman whose humerus is 35 cm long is ??? cm

jimthompson5910 (jim_thompson5910):

H(x) = 2.75x + 71.48 H(35) = 2.75(35) + 71.48 H(35) = 96.25+ 71.48 H(35) = 167.73 So her predicted height is about 167.73 cm

OpenStudy (anonymous):

You essentially substitute "x" with 35, making it F(35)=2.75(35)+71.48 Then: F(35)=96.25+71.48 Which makes F(35)=167.73